OpenStudy (konradzuse):

derivative of 5e^(3(x^2))

OpenStudy (konradzuse):

\[5e^{{3x}^2}\]

OpenStudy (konradzuse):

so I guess to start it off we get rid of the 5 and put it to the left so we worry about only e^blah.

OpenStudy (konradzuse):

u = 3x^2 du =9x So it's going ot be e^u = e^u * u'

OpenStudy (konradzuse):

so that should leave us with 5 * 9x * e^((3x)^2)?

OpenStudy (tennistar):

I cant help sorry not that advanced

OpenStudy (konradzuse):

np ty

OpenStudy (konradzuse):

http://www.wolframalpha.com/input/?i=derivative+of+5e%5E%28%283x%29%5E%282%29%29 This is what wolfman is saying, but I'm not sure if I get exactly how it's happening.

OpenStudy (konradzuse):

I guess we do the (3x)^2 first to get the 9x, but I would also think the derivative of 3 would be needed?

OpenStudy (anonymous):

\[d[5e^{3x^{2}}]/dx= 5e^{3x^{2}}*6x=30x5e^{3x^{2}}\]

OpenStudy (anonymous):

just chain rule

OpenStudy (konradzuse):

That's what I thought so what exactly is wolfman doing?

OpenStudy (anonymous):

idk

OpenStudy (konradzuse):

I guess I see how they are starting off.

OpenStudy (konradzuse):

5e^9x^2

OpenStudy (konradzuse):

18x * 5e^9x^2

OpenStudy (konradzuse):

okay that makes sense now...

OpenStudy (anonymous):

ah sry (3x)^2 or 3x^2

OpenStudy (anonymous):

?

OpenStudy (konradzuse):

(3x)^2

OpenStudy (konradzuse):

I confused myself with that too.

OpenStudy (anonymous):

oh, then is different...

OpenStudy (konradzuse):

actually....

OpenStudy (konradzuse):

I guess it is just 3x^2

OpenStudy (konradzuse):

Im sorry I suck lol :p

OpenStudy (anonymous):

\[d[5e ^{9x ^{2}}]/dx=90x5e ^{9x ^{2}}\]

OpenStudy (anonymous):

sry 5 is a typo

OpenStudy (anonymous):

90xe^9x^2